OpenStudy (anonymous):

if f(x) = 2/x then what is [f(x+h)-f(x)] / h

zepdrix (zepdrix):

Oh the reciprocal thing? Ooo this one is a little stinker to simplify :) lol

OpenStudy (anonymous):

i will attempt

zepdrix (zepdrix):

k

OpenStudy (anonymous):

so for f(x+h) does that mean 2/x+h ?

OpenStudy (anonymous):

and for -f(x) : would that be -2/x

zepdrix (zepdrix):

2/(x+h), yes the brackets are kind of important, otherwise it looks like this \(\Large\rm \frac{2}{x}+h\)

zepdrix (zepdrix):

mhm :d

OpenStudy (anonymous):

lol ok you're right I forgot about that.

OpenStudy (anonymous):

ok so my whole equation will be {[2/(x+h)] - (2/x)} / h

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

\[\Large\rm f(x)=\frac{2}{x}\]\[\Large\rm f(x+h)=\frac{2}{x+h}\]\[\Large\rm \frac{\frac{2}{x+h}-\frac{2}{x}}{h}\] Mmm yes, looks good.

OpenStudy (anonymous):

ok I remember seeing something like this before.

OpenStudy (anonymous):

would I times the equation by h/1?

OpenStudy (anonymous):

that way canceling out the bottom h?

zepdrix (zepdrix):

So we have a couple of fractions in our large numerator. How do we combine fractions? We can't do the subtraction right now, since they don't have the same denominator.

zepdrix (zepdrix):

h/1? Naw we can't just multiply by h/1 willy nilly. Gotta multiply by h/h to keep the equation balanced. But we don't actually want to do that here :d

OpenStudy (anonymous):

oh.

OpenStudy (anonymous):

oh ok. duh h/1 does not equal 1 anyways lol. (wish it were that easy) ok, so.... I think that I can start by taking care of the top two equation

zepdrix (zepdrix):

Yes, do that. They should simplify down into something nice.

OpenStudy (anonymous):

so I have [ 2/(x+h)] and [2/x] I believe their common denominator is x2+xh?

OpenStudy (anonymous):

can i multiply both equations by that?

OpenStudy (anonymous):

so i now have for my numerator(if i am able to multiply both by x^2+xh ) 2x-2x+2h

zepdrix (zepdrix):

Ummm I wouldn't multiply out the denominator like that. It's going to make it harder to see what you actually need in each fraction.

zepdrix (zepdrix):

The first fraction has a factor of \(\Large\rm (x+h)\) while the second fraction has a factor of \(\Large\rm x\) So our common denominator will be \(\Large\rm x(x+h)\), yes?

OpenStudy (anonymous):

oh. ok. i just lost myself

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x^2 +xh

zepdrix (zepdrix):

Nooo, stop multiplying it >.< lol

zepdrix (zepdrix):

Notice that both fractions don't need the entire thing. The first fraction only needs the \(\Large\rm x\) factor, and the second fraction only needs the \(\Large\rm (x+h)\) factor.

OpenStudy (anonymous):

hahahaha ok sorry i have adhd so i get too excited jhahjaha

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so would i have to multiply 2/(x+h) by x ?

OpenStudy (anonymous):

and the second equation 2/x by (x+h) ?

OpenStudy (anonymous):

or am I going in the completely wrong direction?

zepdrix (zepdrix):

Yes good :) \[\Large\rm \color{royalblue}{\frac{x}{x}}\cdot\frac{2}{(x+h)}\]

zepdrix (zepdrix):

\[\Large\rm \color{orangered}{\frac{(x+h)}{(x+h)}}\cdot\frac{2}{x}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so now i have [(2x)/(x^2+xh)] - [(2x+2h)/(2+xh) ]

zepdrix (zepdrix):

Stop multiplying the bottomssssss >:U AHHHHHHHHHHHHHHHHHHHHH

zepdrix (zepdrix):

:3

OpenStudy (anonymous):

oh my goodness i though i was supposed too hahahahaha arent you glad you only have to deal with me tonight ;)

OpenStudy (anonymous):

ok so now what?

zepdrix (zepdrix):

\[\Large\rm \frac{\dfrac{2}{x+h}-\dfrac{2}{x}}{h}\quad=\quad\frac{\dfrac{2x}{x(x+h)}-\dfrac{2x+2h}{x(x+h)}}{h}\]

OpenStudy (anonymous):

oooooooooooooohhhhhhhhhhhhhhhhhhhhhhhh

OpenStudy (anonymous):

thats what you wanted me to do :) ok I promise I am a really smart student, Its just late.

zepdrix (zepdrix):

:3

OpenStudy (anonymous):

ok back to seriousness.

zepdrix (zepdrix):

They have the same denominator, s we can combine the fractions and perform the subtraction if we have any like-terms.

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

\[\Large\rm \frac{\dfrac{2x-2x-2h}{x(x+h)}}{h}\]

zepdrix (zepdrix):

Do you understand why we're subtracting the 2h and not adding it?

OpenStudy (anonymous):

yes becuase the 2x+2h is one term and the negative needs to be distributed to both the 2x and 2h

zepdrix (zepdrix):

k good :)

OpenStudy (anonymous):

ok so I AM still confused on what to do next. and I dont want to jump and try multiplying becuase I am sure I will get it wrong lol

zepdrix (zepdrix):

It looks like our numerator has like-terms. Let's combine the x's.

OpenStudy (anonymous):

yes -2h is the only numerator left

zepdrix (zepdrix):

\[\Large\rm \frac{\frac{-2h}{x(x+h)}}{h}\]Ok cool.

zepdrix (zepdrix):

This next step umm....... Ooo boy I'm at a loss for words what to call this....

zepdrix (zepdrix):

Example: If I have a half, \[\Large\rm \frac{1}{2}\]and I divide that by 3,\[\Large\rm \frac{\left(\frac{1}{2}\right)}{3}\]Realize that there is an invisible 1 here,\[\Large\rm \frac{\left(\frac{1}{2}\right)\cdot1}{3}\]So I can rewrite this as:\[\Large\rm \left(\frac{1}{2}\right)\cdot \frac{1}{3}\]

zepdrix (zepdrix):

Can* you make sense of that crazy fraction math I just did? We're going to apply that to our problem.

OpenStudy (anonymous):

yes i was actually already trying to do that, but I didnt want to get in trouble "jumping ahead " :)

zepdrix (zepdrix):

\[\Large\rm \frac{\left(\frac{-2h}{x(x+h)}\right)}{h}\]lol :P

OpenStudy (anonymous):

{ (-2h) / [x(x+h)] } * [1/h]

OpenStudy (anonymous):

do i need to find another common denominator? or do i cross multiply?

zepdrix (zepdrix):

you -_- you people and your cross multiplication..... sigh...

zepdrix (zepdrix):

We're multiplying fractions so: No, we don't need a common denominator, No, we don't cross multiply.

OpenStudy (anonymous):

i dont know wha * 1/h means lol. looks like a multiply the first equation by 1/h ok so i am wrong, and i am wrong daaarrrnnnnn i love math and i though it was so easy, maybe I am just overthinking things

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

oh so do i just multiply straight across?

zepdrix (zepdrix):

Yes :)

OpenStudy (anonymous):

yyyyayayayayayayayayaya :)

zepdrix (zepdrix):

And again, don't distribute the h to each term in the brackets.

OpenStudy (anonymous):

middle school math = so simple :)

zepdrix (zepdrix):

Just leave it in front :d

OpenStudy (anonymous):

in the front? ok :)

zepdrix (zepdrix):

\[\Large\rm \frac{\left(\frac{-2h}{x(x+h)}\right)}{h}=\left(\frac{-2h}{x(x+h)}\right)\cdot \frac{1}{h}\]

OpenStudy (anonymous):

yes so now I have (-2)/[x(x+h)] because the h on the bottom and top cancel each other out?

zepdrix (zepdrix):

yay :)

zepdrix (zepdrix):

So from this point, you can either leave it like that as your final answer (which is what I would do). Or if you INSIST, then you can multiply out your denominator at this point :d

OpenStudy (anonymous):

hahaha ill leave AS IS. no more multiplying, it keeps getting me in trouble lol

OpenStudy (anonymous):

thank you. i will ATTEMPT the rest of the problems on my own :)

zepdrix (zepdrix):

cool c: gj